# V2X-DGW: Domain Generalization for Multi-agent Perception under Adverse Weather Conditions

### v3

@@ 3.3 Adaptive Weather Augmentation
The AWA process is dynamically executed in real-time for each training iteration. This procedure generates augmented point cloud data, which are exclusively utilized for the purpose of simulating the unseen adverse weather conditions, so as to improve the model generalization capability.
3.4 Trust-region Weather-invariant Alignment (TWA)
((a))
@@ 3.4 Trust-region Weather-invariant Alignment (TWA)
$\mathbf{T}(h,w)=\begin{cases}1&\text{if }\exists c\in\{1,\ldots,C\}:\mathbf{I}^{s}(c,h,w)\neq 0\land\mathbf{I}^{r}(c,h,w)\neq 0\\ 0&\text{otherwise}.\end{cases}$ (7)
@@
3.5 Agent-aware Contrastive Alignment (ACA)
We propose Agent-aware Contrastive Alignment (ACA), as contrastive learning [13] based regularization, to relieve accumulated damages in multi-agent system during the adverse weather conditions.
@@ 3.5 Agent-aware Contrastive Alignment (ACA)
$\mathcal{L}_{ACA_{a}}=-\frac{1}{B_{a}}\sum_{i=1}^{B_{a}}\sum_{p\in P(i)}\log\frac{\exp\left(\mathbf{F}_{i}^{s}\cdot\mathbf{F}_{p}^{s}/{\tau}\right)+\exp\left(\mathbf{F}_{i}^{s}\cdot\mathbf{F}_{p}^{a}/{\tau}\right)+\exp\left(\mathbf{F}_{i}^{a}\cdot\mathbf{F}_{p}^{a}/{\tau}\right)}{\sum_{j=1}^{B_{a}}\exp\left(\mathbf{F}_{i}^{s}\cdot\mathbf{F}_{j}^{s}/{\tau}\right)+\sum_{k=1}^{B_{a}}\exp\left(\mathbf{F}_{i}^{a}\cdot\mathbf{F}_{k}^{a}/{\tau}\right)},$ (10)
@@
$\mathcal{L}_{ACA_{g}}=-\frac{1}{B_{g}}\sum_{i=1}^{B_{g}}\log\frac{\exp\left(\hat{\mathbf{F}}_{i}^{s}\cdot\hat{\mathbf{F}}_{i}^{a}/{\tau}\right)}{\sum_{j=1}^{B_{g}}\exp\left(\hat{\mathbf{F}}_{i}^{s}\cdot\hat{\mathbf{F}}_{j}^{s}/{\tau}\right)+\sum_{k=1}^{B_{g}}\exp\left(\hat{\mathbf{F}}_{i}^{a}\cdot\hat{\mathbf{F}}_{k}^{a}/{\tau}\right)},$ (11)
@@
3.6 Overall Loss Function
4 Experiments
4.1 Experiment Setup
Dataset. We utilize two popular multi-agent perception datasets for the domain generalization task for point cloud-based 3D object detection. OPV2V [48] is a comprehensive simulated dataset designed for V2V cooperative perception tasks, collected through the integration of the CARLA [7] and OpenCDA [46] platforms. It encompasses 73 varied scenes with CAVs. The dataset includes 6,764 frames for training and 2,719 frames for testing. V2XSet [47] is an expansive simulated dataset for V2X cooperative perception task, is also assembled utilizing CARLA. It features LiDAR data from a variety of autonomous vehicles and roadside infrastructures, synchronized within identical scenarios. The dataset is organized into training and testing subsets, containing 6,694 and 2,833 frames respectively. Here, we simulate adverse weather conditions only on the testing sets of these two datasets to generate OPV2V-w and V2XSet-w datasets for our experiments, as detailed in Section 3.1.
Compared Methods. To analyze the effect of adverse weather conditions on the perception performance of multi-agent system, two Domain Generalization (DG) methods, i.e., IBN-Net [34], and MLDG [18], and three Domain Adaptation (DA) methods, i.e., GRL [8], AdvGRL [24], and S2R-AFA [22] are implemented to be evaluated.
[…]
Experimental Settings. This paper aims to enhance the perception performance of multi-agent systems via domain generalization technology under unseen adverse weather scenarios. We evaluate the detection performance across four weather conditions: clean, fog, rain and snow. All methods in the experiments employ the same cooperative perception method AttFuse [48] (as FFN), with the PointPillar based Encoder for fair comparison. We denote the AttFuse method without DA or DG as our Baseline, which is only trained on source data (clean weather) without AWA. For implementation of our V2X-DGW method, $\phi_{L}$ and $\phi_{U}$ are set to 0.5 and 0.8, $\alpha_{1}$ and $\alpha_{2}$ are set to 0.1 and 1, $\beta_{1}$ and $\beta_{2}$ are set to 0.01 and 0.01, $\tau$ is set to 0.07. We assess all methods under two key settings: 1) OPV2V2ALL. All DG methods including ours are trained on the OPV2V training set only. For DA methods, the labeled OPV2V training set serves as the source domain, while the unlabeled OPV2V training set added fog by Physics-based Simulation is utilized as the target domain during training. 2) V2XSet2ALL. In this setting, all DG methods including ours are trained on the V2XSet training set only. For DA methods, the labeled V2XSet training set is as the source domain, and the unlabeled V2XSet training set added fog by Physics-based Simulation is as the target domain during the training. For the above both settings, these methods are all tested/evaluated on the OPV2V-w and V2XSet-w with fog, rain, snow weather conditions.
Table 1: 3D detection performance on OPV2V2ALL setting. We show Average Precision (AP) at IoU=0.5, 0.7 on four weather conditions, respectively.
@@ 4.1 Experiment Setup
4.2 Quantitative Evaluation
@@
Performance in OPV2V2ALL. Table 1 shows the performance comparison on OPV2V2ALL setting, where all methods are evaluated on OPV2V, OPV2V-w and V2XSet-w testing sets, respectively. Without the effect of the adverse weathers, the Baseline without any DA or GD methods can gain a outstanding performance on OPV2V clean testing set. However when it is deployed on OPV2V-W testing sets, it drops $27.59\%$ / $32.52\%$ for AP@0.5/0.7 on the snow weather of OPV2V-W. It indicates that the cooperative perception system has significantly detrimental effects on performance under adverse weather conditions, which also underscores the importance of generalization capabilities. After applying the DA and GA methods, only a few have shown improved performance on the unseen weather domains.This could be attributed to the differences in domain gaps between 2D images and point clouds, as well as the ambiguity of appearance features in the LiDAR domain. Consequently, the performance enhancement achieved through learning appearance-invariant features may not be sufficient. While our method consistently outperforms comparative benchmarks in both clean and adverse weather conditions, as evidenced by the results presented in Table 1. For examples, our proposed method is improved by $5.29\%$/$11.42\%$ for for AP@0.5/0.7 on the snow weather of OPV2V-w. It underscores the generalization capabilities of our proposed method for cooperative perception across a variety of adverse weather conditions.
Table 2: 3D detection performance comparison with V2XSet2ALL setting. We show Average Precision (AP) at IoU=0.5, 0.7 on clean, fog, rain, snow weather conditions, respectively.
V2XSet OPV2V-w V2XSet-w
clean fog rain snow fog rain snow
Baseline 81.62 70.68 67.16 58.98 70.97 59.85 59.53 47.70 64.16 57.09 66.11 56.60 67.27 57.61
IBN-Net [34] 81.64 71.13 65.29 57.46 67.29 56.93 54.52 43.73 60.53 52.83 61.64 52.15 63.77 54.25
[…]
Ours 84.14 74.53 67.53 59.56 72.63 62.75 61.78 50.57 65.17 57.99 69.94 61.02 68.61 59.53
GRL [8] 82.67 72.80 67.17 58.87 69.95 59.11 58.76 47.88 62.90 56.18 64.45 55.48 66.38 56.76
AdvGRL [24] 82.64 72.09 67.26 58.49 67.50 56.35 59.46 48.59 63.87 55.63 63.35 52.81 64.43 53.82
S2R-AFA [22] 80.71 70.32 66.95 58.91 69.52 56.67 59.08 45.31 62.73 55.51 63.14 51.44 64.50 52.77
Performance in V2Vset2ALL. The LiDAR-based 3D object detection results on the V2VSet2ALL setting are presented in Table 2. Among all DA and GA methods, our proposed GA method achieves the best performance by eliminating the domain gap generated by the adverse weather conditions in the OPV2V-w and V2XSet-w testing sets. The outcomes under adverse weather conditions are akin to those observed when trained on the OPV2V2ALL setting. This similarity is attributable to both V2XSet and OPV2V being collected within the CARLA simulator [7], thereby mitigating extraneous influences and focusing on the impact of weather variations on multi-agent perception systems. Our method also achieve the best performance under adverse weather conditions.
Sensitivity to agent numbers. In a multi-agent perception system under adverse weather, the ego agent might be particularly vulnerable to the accumulated damages from other agents. Fig. 4 illustrates the impact of cooperative perception performance with varying numbers of agents under adverse weather. Intuitively, the cooperative perception performance should be improved with an increasing number of nearby agents/CAVs [48]. However, most comparison methods such as Baseline, GRL, AdvGRL, S2R-AFA, IBN-Net sometimes have decreasing cooperative detection performance with increasing agent numbers under some adverse weather scenarios. MLDG’s performance is sometimes keeping almost unchanged with the increasing agent numbers. This unreasonable phenomena might be due to their sensitivity to the accumulated damages from more agents in adverse weather. Differently, our V2X-DGW has continually improved cooperative detection performance with increasing agent numbers.
((a))
Figure 4: Sensitivity to agent numbers. Detection (AP) with different agent/CAV numbers on OPV2V-w under adverse weather conditions.
Accumulated damages in V2X. As depicted in (a) and (b) of Fig. 5, a single CAV can detect the highlighted vehicle under rainy weather conditions, whereas a multi-agent perception system with two CAVs even miss that vehicle within their perception range by the Baseline method. This type of accumulated damages caused by adverse weather can impair cooperative perception performance in real-world scenarios. However, our proposed V2X-DGW method demonstrates enhanced robustness against such cumulative damages under adverse weather.
((a))
((b))
((c))
((d))
Figure 5: Illustration of accumulated damages in the multi-agent system under adverse weather. One frame from V2XSet-w rain is shown as example here. Green and red 3D bounding boxes represent the ground truth and prediction respectively.
3D detection visualization. Figure 6 presents the 3D detection visualization on OPV2V-w fog subset of the Baseline, S2R-AFA [22], and our proposed method. In clean weather, nearly all methods generate bounding boxes with high accuracy, closely aligning with the ground truths. However, when it is generalized to adverse weather conditions, both the baseline and DA methods exhibit detection errors. More critically, our method succeeds in identifying a greater number of dynamic objects, as evidenced by a higher match rate with ground-truth. This underscores the superior generalization capabilities of our proposed method.
@@
Representation visualization. We visualize the agent-level and group-level representation in V2X perception system using T-SNE [30]. As shown in Fig. 7, samples within the representation space resemble curves, attributable to the OPV2V-w consisting of continuous temporal frames. At both the agent-level and group-level, our proposed V2X-DGW method demonstrates a more cohesive clustering of samples across the four different weather conditions, as indicated by the tighter grouping of differently colored curves. This phenomenon is because of the implementation by our two alignments TWA and ACA. In other words, the generalizable weather-invariant features are learned by our V2X-DGW.
@@
((b))
((c))
((d))
Figure 7: Representation visualization with T-SNE [30]. We show two-type features in multi-agent system: Agent-Level (AL) and Group-Level (GL). Every point represents a data sample, and different color represents its different weather condition.
Ablation studies. To validate the effectiveness of our proposed components, we provide the quantitative comparison on detection result under clean, fog, rain, snow weather conditions in Table 3. Each component contributes to the model’s performance enhancement across all four weather conditions, validating the rationale behind the design of AWA, TWA, and ACA.
Table 3: Ablation study. Effects of the AWA, TWA, ACA components. The first row is the result by Baseline.
@@
AP@0.5 AP@0.7 AP@0.5 AP@0.7 AP@0.5 AP@0.7 AP@0.5 AP@0.7
88.10 79.48 67.54 59.39 71.17 58.33 60.51 46.96
87.03 79.31 69.42 62.58 74.16 66.50 64.14 56.73
88.88 81.15 70.83 63.71 74.36 67.10 64.65 57.21
88.66 80.80 70.95 63.74 74.90 67.42 65.06 57.63
89.23 81.15 71.35 64.34 75.30 67.97 65.80 58.38
5 Conclusion
In this paper, we propose a domain generalization approach for LiDAR-based 3D object detection on multi-agent perception system under adverse weather conditions. To investigate weather impact on multi-agent perception system, the effect of three common adverse weather conditions are simulated on the two publicized V2V/V2X datasets, and to obtain two adverse weather datasets: OPV2V-w and V2XSet-w. To address the range reduction, point cloud degradation and accumulated damages in adverse weather, we first introduce the Adaptive Weather Augmentation (AWA) to simulate the unseen adverse weather domain, and then propose two alignments for generalizable representation learning: Trust-region Weather-invariant Alignment (TWA) and Agent-aware Contrastive Alignment (ACA). Extensive experiments show that our method achieved significantly improved performance in the unseen adverse weather conditions.

### v5

@@ III-C Adaptive Weather Augmentation
The AWA process is illustrated in Fig. 3, which is dynamically executed in real-time during each training iteration. This procedure generates augmented point cloud data, which is exclusively utilized for the purpose of simulating the unseen adverse weather conditions, so as to improve the model generalization capability.
III-D Trust-region Weather-invariant Alignment (TWA)
@@
$\mathbf{T}(h,w)=\begin{cases}1&\text{if }\exists c\in\{1,\ldots,C\}:\mathbf{I}^{s}(c,h,w)\neq 0\land\\ &\mathbf{I}^{r}(c,h,w)\neq 0\\ 0&\text{otherwise}.\end{cases}$ (6)
@@
III-E Agent-aware Contrastive Alignment (ACA)
We propose Agent-aware Contrastive Alignment (ACA) as contrastive learning [36] based regularization to enhance the model robustness so as to relieve the accumulated damages on multi-agent perception under the adverse weather conditions.
@@ III-E Agent-aware Contrastive Alignment (ACA)
$\mathcal{L}_{ACA_{a}}=-\frac{1}{B_{a}}\sum_{i=1}^{B_{a}}\sum_{p\in P(i)}\Bigg{\{}\log[\exp\left(\mathbf{F}_{i}^{s}\cdot\mathbf{F}_{p}^{s}/{\tau}\right)\\ +\exp\left(\mathbf{F}_{i}^{s}\cdot\mathbf{F}_{p}^{a}/{\tau}\right)+\exp\left(\mathbf{F}_{i}^{a}\cdot\mathbf{F}_{p}^{a}/{\tau}\right)\Big{]}\\ -\log[\sum_{j=1}^{B_{a}}\exp\left(\mathbf{F}_{i}^{s}\cdot\mathbf{F}_{j}^{s}/{\tau}\right)+\sum_{k=1}^{B_{a}}\exp\left(\mathbf{F}_{i}^{a}\cdot\mathbf{F}_{k}^{a}/{\tau}\right)\Big{]}\Bigg{\}},$ (9)
@@
$\mathcal{L}_{ACA_{g}}=-\frac{1}{B_{g}}\sum_{i=1}^{B_{g}}\Biggl{\{}\log[\exp\left(\hat{\mathbf{F}}_{i}^{s}\cdot\hat{\mathbf{F}}_{i}^{a}/{\tau}\right)\Biggr{]}\\ -\log[\sum_{j=1}^{B_{g}}\exp\left(\hat{\mathbf{F}}_{i}^{s}\cdot\hat{\mathbf{F}}_{j}^{s}/{\tau}\right)+\sum_{k=1}^{B_{g}}\exp\left(\hat{\mathbf{F}}_{i}^{a}\cdot\hat{\mathbf{F}}_{k}^{a}/{\tau}\right)\Biggr{]}\Biggr{\}},$ (10)
@@
TABLE I: 3D detection performance from clean weather to adverse weather. The training process utilizes only the source data (clean weather). During inference, the trained model can be applied to clean, foggy, rainy, and snowy weathers.
@@
III-F Overall Loss Function
IV Experiments
IV-A Experiment Setup
Dataset. We utilize two popular multi-agent perception datasets. OPV2V [4] is a comprehensive simulated dataset designed for V2V cooperative perception tasks, collected through the integration of the CARLA [35] and OpenCDA [42] platforms. It encompasses 73 varied scenes with CAVs. The dataset includes 6,764 frames for training and 2,719 frames for testing. V2XSet [8] is an expansive simulated dataset for V2X cooperative perception task, is also assembled utilizing CARLA. The dataset is organized into training and testing subsets, containing 6,694 and 2,833 frames respectively. We simulate adverse weather conditions only on their testing sets to generate OPV2V-w and V2XSet-w datasets for our experiments, as detailed in Section III-A.
Compared Methods. To analyze the effect of adverse weather conditions on the perception performance of multi-agent system, two Domain Generalization (DG) methods, i.e., IBN-Net [37], and MLDG [38], and three Domain Adaptation (DA) methods, i.e., GRL [39], AdvGRL [40], and S2R-AFA [5] are implemented to be evaluated. For the MLDG [38] setting, we employ two different strategies: randomly splitting each batch into meta-train and meta-test sets, denoted as MLDGa, and splitting data by maximizing feature distances, denoted as MLDGb. As for the domain adaptation methods, we follow the settings in [5].
Experimental Settings. We evaluate the detection performance across four weather conditions: clean, fog, rain and snow. All methods in the experiments employ the same cooperative perception method AttFuse [4] (as FFN), with the PointPillar based Encoder for fair comparison. We denote the AttFuse method without DA or DG as our Baseline, which is only trained on source data (clean weather). For implementation of our V2X-DGW method, $\phi_{L}$ and $\phi_{U}$ are set to 0.5 and 0.8, $\alpha_{1}$ and $\alpha_{2}$ are set to 0.1 and 1, $\beta_{1}$ and $\beta_{2}$ are set to 0.01 and 0.01, $\tau$ is set to 0.07. All DG methods are trained on the OPV2V training set (clean weather) only. For DA methods, the labeled OPV2V training set serves as the source domain, while the unlabeled OPV2V training set added with fog by physics-based simulation is utilized as the target domain during training. These methods are all tested/evaluated on the OPV2V-w and V2XSet-w with fog, rain, snow weather conditions.
IV-B Quantitative Evaluation
@@
3D Detection Performance. Table I shows the experiment performance. Without the effect of the adverse weathers, the Baseline without any DA or DG methods can gain a outstanding performance on OPV2V clean testing set. However, when it is deployed on OPV2V-w testing sets, it drops $27.59\%$ / $32.52\%$ for AP@0.5/0.7 on the snow weather of OPV2V-w. It indicates that the cooperative perception system has significantly detrimental effects on performance under adverse weather conditions, which also underscores the importance of generalization capabilities. After applying the DA and DG methods, only a few have shown improved performance on the unseen weather domains. This could be attributed to the differences in domain gaps between 2D images and point clouds, as well as the ambiguity of appearance features in the LiDAR domain. Consequently, the performance enhancement achieved through learning appearance-invariant features may not be sufficient. While our method consistently outperforms comparative benchmarks in both clean and adverse weather conditions. It underscores the generalization capabilities of our proposed method for cooperative perception across a variety of adverse weather conditions.
@@
3D Detection Visualization. The Fig. 5 presents the 3D detection visualization on OPV2V-w fog subset of the Baseline, S2R-AFA [5], and our proposed method. When generalizing a cooperative perception model trained in clear weather to adverse weather conditions, both the baseline and DA methods exhibit detection errors, highlighting the detrimental impact of harsh weather on cooperative perception. More critically, our method succeeds in identifying a greater number of dynamic objects, as evidenced by a higher match rate with ground-truth. This underscores the superior generalization capabilities of our proposed method.
@@
Figure 6: Sensitivity to agent numbers. Other methods’ decreasing detection performance with more agent numbers on OPV2V-w indicates the accumulated damages under adverse weather, while our method is more robust.
Sensitivity to Agent Numbers. The Fig. 6 illustrates the impact of cooperative perception performance with varying numbers of agents under rain and snow weather. Intuitively, the cooperative perception performance should be improved with an increasing number of nearby agents [4]. However, most comparison methods sometimes have decreasing cooperative detection performance with increasing agent numbers under adverse weather scenarios, which is because of their sensitivity to the accumulated damages from more agents in adverse weather. Differently, our V2X-DGW could overcome the challenge of accumulated damages.
Ablation Studies. To validate the effectiveness of our proposed components, we provide the quantitative comparison on detection result under clean, fog, rain, snow weather conditions in Table II. Each component contributes to the model’s performance enhancement across all four weather conditions, validating the rationale behind the design of AWA, TWA, and ACA.
TABLE II: Ablation study. Effects of the AWA, TWA, ACA components. The first row is the result by Baseline.
@@
AP@0.5/0.7 AP@0.5/0.7 AP@0.5/0.7 AP@0.5/0.7
88.10/79.48 67.54/59.39 71.17/58.33 60.51/46.96
87.03/79.31 69.42/62.58 74.16/66.50 64.14/56.73
88.88/81.15 70.83/63.71 74.36/67.10 64.65/57.21
88.66/80.80 70.95/63.74 74.90/67.42 65.06/57.63
89.23/81.15 71.35/64.34 75.30/67.97 65.80/58.38
V Conclusions
In this paper, we propose a domain generalization approach for LiDAR-based 3D object detection on the multi-agent perception systems under the adverse weather conditions (Fog, Rain, Snow). To investigate the weather impact on multi-agent perception, the effect of three common adverse weather conditions are simulated on the two publicized V2X datasets, and to obtain two new V2X based adverse-weather multi-agent perception datasets: OPV2V-w and V2XSet-w. To address the range reduction, point cloud degradation and accumulated damages in adverse weather, we first introduce the Adaptive Weather Augmentation (AWA) to simulate the unseen adverse weather domain, and then propose two alignments for generalizable representation learning: Trust-region Weather-invariant Alignment (TWA) and Agent-aware Contrastive Alignment (ACA). Extensive experiments show that our method achieved outstanding 3D detection performance in the unseen adverse weather conditions.